OpenStudy (kirbykirby):

For any n, 33 divides (n^101-n). Prove this is true.

OpenStudy (kirbykirby):

There's a solution given and it gives the hint that you want to prove "n\[n^{100}\approx n(mod 33)\]. I'm using the \[\approx\] for "congruent to". There's a solution which I partially understand. The only step confusing me is at the beginning is when they say: From Fermat's theorem, \[n ^{2}\approx1(mod 3) => n ^{3}\approx n(mod 3)\] and \[n ^{10}\approx1(mod 11) => n ^{11}\approx n(mod 3)\] How would this be true for EVERY n (as stated initially) since Fermat's theorem only works if \[p\] doesn't divide \[a\] (where p is prime) in \[a ^{p-1}\approx1(\mod p)\]

OpenStudy (kirbykirby):

Oops I made a mistke, it should be n^11 = n (mod 11) ... not mod 3!

OpenStudy (anonymous):

kirby p is the exponent, not n

OpenStudy (anonymous):

n can be anything at all

OpenStudy (kirbykirby):

but wouldn't it have to be in this case that 3 doesn't divide n and that 11 doesn't divide n ? Wouldn't this not work if n was a multiple of 3 or 11 respectively

OpenStudy (anonymous):

oooh i see

OpenStudy (anonymous):

but it is trivially true if p divides n

OpenStudy (anonymous):

if p divides n then n^p is a multiple of p, but is certainly congruent to n mod p because n mod p is 0

OpenStudy (anonymous):

for example 3 divides 9 and 3^9=27 congruent to 9 mod p congruent to 0 mod p

OpenStudy (anonymous):

i have never seen fermat's little theorem stated that n and p had to be relatively prime. just looked it up on wiki and it doesn't say that either

OpenStudy (kirbykirby):

hm weird, my book clearly stated it o_O And they said the same for the more general Euler's theorem. But I guess I'll take your word for it :)

OpenStudy (anonymous):

hold on. if p divides n, then n is congruent to n mod p because n mod p is 0

OpenStudy (anonymous):

There are 2 formulations of FLT 1) a^(p-1) = 1 mod p with gcd(a,p)=1 OR 2) a^p = a mod p if p is a prime and a any integer.

OpenStudy (anonymous):

@kirby you are trying to prove that for any n, 33 divides \[n^{101}-n\] if 33 already divides n you have nothing to prove because if it divides n it certainly divides \[n^{101}-n\]

OpenStudy (anonymous):

so you have no worries in that case. it is true before you start

OpenStudy (kirbykirby):

Ohh! Haha oh my , that should've been obvious from the start :( Haha thanks a lot :)

OpenStudy (kirbykirby):

Oh but wait! I think I'm getting confused! How do we know 33 divides n from the start?

OpenStudy (kirbykirby):

Ohh ok I see from estudier's post that the condition that (a, p)=1 is "removed" if you state it as a^p = a (mod p)... That was not really said in my book, or I didn't catch that.. Ok now it makes sense :)